OpenStudy (anonymous):

In the isosceles right triangles ABC and MNP, AB = BC = 12 and MN = NP = 24. Which is the sum of the hypotenuses of these triangles? 16 square root 2 18 square root 2 32 square root 2 36 square root 2 i got the 2cd one..

OpenStudy (anonymous):

Using the pythagorean theorem we can get the hypotenuses of the triangles: \[h_{1} = \sqrt{12^{2}+12^{2}} = \sqrt{2(12)^{2}} = 12\sqrt{2}\] Likewise the second hypotenuse will be \[24\sqrt{2}\] adding them gives \[36\sqrt{2}\]

OpenStudy (blacksteel):

Well, by the Pythagorean Theorem, the hypotenuse of the first triangle is \[\sqrt(12^2 + 12^2) = \sqrt(2*12^2) = 12*\sqrt(2)\] and the hypotenuse of the second one is \[\sqrt(24^2 + 24^2) = \sqrt(2*24^2) = 24*\sqrt(2)\] Then the sum of these is (24+12) square roots of 2, or \[36*\sqrt(2)\]

myininaya (myininaya):

gj joe

OpenStudy (anonymous):

thank you :)

OpenStudy (blacksteel):

Curses, Joe beat me to it!

OpenStudy (anonymous):

gj Blacksteel :)

myininaya (myininaya):

gj blacky!

OpenStudy (anonymous):

its not a race! (although sometimes it feels like it is <.<) lol

myininaya (myininaya):

no no it is a race

myininaya (myininaya):

lol